OpenStudy (anonymous):

Simplify the following expression. 3/4 + 1/60 * 23/60 23/30 1/16 41/60

OpenStudy (anonymous):

All you need to do is make sure you have common denominator.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

Yes, you need an LCD. So you have: \[\dfrac{ 3 }{4 }+\dfrac{ 1 }{60 }\] You have rewrite with equivalent fractions that have the same denominator. So you want the LCD - which is the LCM (least common multiple) of 4 and 60. Since 60 is a multiple of 4 (60=15*4) you have 60 as your LCD. So you need to muliply the first fraction ("build up") so that it has a den'r of 60: e.g. \(\dfrac{ 3 }{4 }\cdot\dfrac{ a }{a }=\dfrac{ 3a }{60 }\) Just figure out what "a" need to be, in order to give you the 60 in the den'r, then multiply by that same thing in the num'r.

OpenStudy (debbieg):

That gives you a fraction that is EQUIVALENT to 3/4, but has a den'r of 60, allowing you to add it to the other fraction.

OpenStudy (anonymous):

23/30